OpenStudy (anonymous):

Write the following equation in the form y=mx+b. 1. 8x-2y=6 2. -y+2x-7=0 .please can someone help me...!!!!

OpenStudy (larseighner):

All you have to do here is solve for y. It will automagically be in the required form. Here is one for starters. 8x- 2y = 6 subtract 8x from both sides -2y = -8x + 6 divide through by -2 y = 4x - 3 That's it. m = 4, b = -2.

OpenStudy (anonymous):

ok.I get the #1 but i don't get the #2.

OpenStudy (anonymous):

-y+2x-7=0 -y=2x-7+0 \[\frac{ -y }{ -1}=\frac{ 2x-7 }{ -1}\] y=-2x+7 m=-2 b=7 Can someone please check my answer is correct.??????

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!}\)

OpenStudy (larseighner):

Well not quite. when you move the 2x - 7 across the = you are actually subtracting them from both sides, so their signs have to change. In this case, it is much easier to add y to both sides. -y+2x-7=0 add y to both sides 2x - 7 = y since equals signs work both ways y= 2x - 7

OpenStudy (anonymous):

Why 2x become a positive and the 7 became negative.?

OpenStudy (larseighner):

This is the original equation (from your question above) -y+2x-7=0 You can add y to both side as I just demonstrated, or you can subtract 2x - 7 from both sides. If you do the latter it goes: -y+2x-7=0 -y+2x-7 - (2x - 7)=0 - (2x - 7) -y = -2x + 7 then divide both sides by -1 (or multiply through by -1) y = 2x - 7